Adam is 5 years younger than Eve. In 1 year, Eve will be 3 times as old as Adam was 4 years ago. Find their ages now.
Thanks for your help!

Answer by mananth(16946)

let eve be x years
Adam is x-5 years
...
In 1 year their ages will be
Eve = x+1 years
Adam = x-5+1 = x-4
..
4 years ago
Adam was x-5-4 = x-9 years.
...
x+1 = 3(x-9)
x+1 = 3x-27
x-2x=-27-1
-x=-28
x= 28 Eve's age now
Adam's age now = 28-4 = 24 years
...
m.ananth@hotmail.ca